OpenStudy (anonymous):

Hi! I need help solving this, but I don't want to solve this one. Can you give me a problem similar to this and help me work through it? What is the solution of the equation? 2p + 21 = 6p – 15 P=___ Do not answer the question above! Help me with a new one!

OpenStudy (anonymous):

p=8

OpenStudy (anonymous):

No, please delete that. Did you not read what I just wrote?! I didn't want the answer!

OpenStudy (superflytyguy):

Oh

OpenStudy (anonymous):

I wanted you to come up with a SIMILAR problem and help me work through it!

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

For pretty much any "solve for the variable" questions, you basically work the order of operations in reverse until your variable is alone and you know what that variable is. \(ax+b=cx+d\) PEMDAS becomes SADMED So, any added/subtracted? Yes. So do the opposite to move things around: \(ax + b - cx=cx+d - cx\) \(ax+b - cx=cx+d - cx\) \(ax- cx+b =d \) \(ax- cx+b-b =d -b\) \(ax- cx =d -b\) \((a- c)x =d -b\) now all my x things are on one side and my other stuff on the other.. but there are still constants multiplied so I would need to divide. \(x =(d -b)/(a- c)\) Now, I can also show it with real numbers...

OpenStudy (triciaal):

here are the steps that will always work isolate the variables to one side of the equation and the numbers to the other. whatever you do to one side of an equation always do to the other to keep the balance. simplify divide each side by the coefficient of the variable

OpenStudy (triciaal):

@CrazyCountryGirl do you want to review these steps to the problem posted before attempting another?

OpenStudy (triciaal):

by the way the answer posted by @prittygirl799 was incorrect.

OpenStudy (anonymous):

@triciaal No. I'd rather do another problems. Can you just replace the numbers and variables and help me work through it. I don't want the answer to that one because it's from a test and so I don't want to cheat. I was just using it as an example

OpenStudy (triciaal):

try this one: -3(3x+15)-(10+x)=35 to use the distributive property just multiply each term inside the parenthesis by what is outside this gets rid of the parenthesis

OpenStudy (e.mccormick):

Just a worked example: Whatever I do to one side of the = I do to the other: Get all the constants to one side: 7x-4=x+32 7x-4+4=x+32+4 7x=x+36 Get all the variables to the other side: 7x=x+36 7x-x=x-x+36 6x=36 Multiply/divide off any coefficents. 6x=36 6x/6=36/6 x=6 And there is my answer.

OpenStudy (e.mccormick):

Now, if you get all of how I did that and why, great! But if not, ask about it. Then we can deal with what you find hard. =)